OpenStudy (chris215):

If f(x) = 9x10 tan-1x, find f '(x).

OpenStudy (chris215):

I got 90x^9tan^-1x + 9x^10 1/1+x^2

HanAkoSolo (jamierox4ev3r):

hmm.. I'm not sure if I'm completely clear on your notation. is this your function? \(\Large f(x)=9x10 tan(-1x)\) or \(\Large f(x)=9x10tan(x)-1x\)

OpenStudy (chris215):

opps sorry its f(x) = 9x^10 tan^-1x

OpenStudy (misty1212):

HI!!

OpenStudy (chris215):

hi :)

OpenStudy (misty1212):

it is a product, you need to use the product rule

OpenStudy (misty1212):

\[\left(fg\right)'=f'g+g'f\] with \[f(x)=9x^{10},f'(x)=90x^8,g(x)=\tan^{-1}(x), g'(x)=\frac{1}{1+x^2}\]

HanAkoSolo (jamierox4ev3r):

There is also an instance of chain rule that you have to take into account when using the product rule for this particular derivation

OpenStudy (misty1212):

oh i didn't even see your answer! it is right !!

OpenStudy (chris215):

oh ok thank you!!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

HanAkoSolo (jamierox4ev3r):

lol @misty1212 , well done breaking down the product rule! I think that was very well explained ^_^